OpenStudy (anonymous):

Solve for x: –4(3x – 2) = 6x + 2 a.1/3 b.6/13 c.-10/13 d.-5/9

OpenStudy (anonymous):

@Cookiemonster32

OpenStudy (anonymous):

–4(3x – 2) = 6x + 2 -12x + 8 = 6x + 2 -12x + 8 = 6x + 2 -6x -6x -6x + 8= 2 - 8 -8 -6x= -6 -6x= -6/-6 = 1 x=1